# Molecule Language Model with Augmented Pairs and Expertise Transfer

### v1

@@ 5.3 Zero-Shot Question and Answering
The molecule is positive w.r.t. a property that is defined as the ability of the body
to be absorbed from the human gastrointestinal system into the bloodstream $\cdots$
@@
This molecule is known to inhibit P-glycoprotein, which is an ABC transporter
protein involved in intestinal absorption, drug metabolism, and brain $\cdots$
@@
This molecule is active w.r.t. Vitamin D receptor. The best pharmacophore
hypothesis contains one hydrogen bond acceptor (A), one hydrogen bond $\cdots$
Table 5: Examples of abstractive and detailed prompts for each dataset.
Figure 3: Hit rate (%) in virtual screening task.
5.4 Molecular Property Prediction
Task Description. In this task, we assess the potential benefits of incorporating external knowledge, i.e., textual descriptions, into the molecule encoder $f_{\text{mol}}$ as done in AMOLE, in enhancing the prediction of molecular properties. We mainly compare to recent graph self-supervised learning (GraphSSL) approaches, i.e., AttrMask Hu et al. (2019), ContextPred Hu et al. (2019), GPT-GNN Hu et al. (2020), InfoGraph Sun et al. (2019), MolCLR Wang et al. (2022), GraphMVP Liu et al. (2021), and Mole-BERT Xia et al. (2022), and MoLMs which represent molecules with graph structure (MoLM w/ Graph). Following previous works Liu et al. (2021, 2023a), we pre-train the molecule encoder $f_{\text{mol}}$ using each of the proposed strategies, and fine-tuning on MoleculeNet benchmark Wu et al. (2018). We provide further details on the dataset and evaluation scheme for the task in Appendix C.4 and D.3, respectively.
Empirical Results. In Table 4, we have following observations: 1) We observe that incorporating external textual descriptions into the pre-training phase uniformly enhances the prediction of molecular properties, as evidenced by improved overall performance (i.e., averaged across all eight tasks). This improvement is credited to the implicit influence of external domain knowledge, i.e., the textual descriptions of molecules, which is proved to be beneficial for property prediction. Liu et al. (2023a). 2) Among the MoLM, AMOLE outperforms baseline methods on six out of eight tasks, demonstrating that the knowledge can be more efficiently transferred to molecular property prediction through our strategy. In summary, our methodology is advantageous not just for cross-modal tasks, but also for tasks within a single modality, demonstrating the versatility of AMOLE across a wide range of downstream applications.
5.5 Zero-Shot Virtual Screening
@@
Task Description. Virtual screening is a computational technique to search large libraries of compounds quickly to identify those structures most likely to have desired properties, emerging as a principal technique in the drug discovery process Mehta et al. (2021). Therefore, in this paper, we propose a novel task named “Zero-shot Virtual Screening,” where we assess the model’s proficiency in virtual screening drugs by supplying a textual description of a desired property. Specifically, our evaluation of the model’s capability in virtual screening is conducted by providing two distinct prompts for each property as shown in Table 5: one brief and abstract, and the other longer and more detailed, offering comprehensive information about the property. For each description, we identify the top 100 molecules nearest to the prompt in the representation space and compute the hit rate to evaluate the model’s performance. We provide further details on each dataset and used prompts for virtual screening in Appendix C.5 and D.4.
@@
Empirical Results. In Figure 3, we have the following observations: 1) While previous studies have shown comparable results on various downstream tasks, their effectiveness in virtual screening leaves room for improvement. Notably, these models often yield results inferior to the mere random selection of molecules (i.e., gray dashed line in Figure 3), and their performance significantly fluctuates based on the specific textual description used. 2) However, AMOLE consistently performs the best in various datasets and different types of descriptions, further demonstrating its ability to learn a more qualified joint space between molecules and text. 3) One interesting observation is that, AMOLE demonstrates notably robust performance compared to the baseline methods, regardless of whether the description is abstract or detailed. This can be attributed to the expertise transfer module, which equips the model with the ability to deduce related information even when only an abstract level of detail is provided. In conclusion, AMOLE can consistently screen the qualified molecules given any textual description, highlighting the adaptability of AMOLE for real-world drug discovery process. We provide additional experimental results on more datasets, how the expertise transfer module affects the model performance with various levels of $\alpha$ in Appendix E.3.
6 Conclusion
In this paper, we propose AMOLE, addressing the two unique challenges in MoLM, i.e., scarcity of molecule-text paired data in both quantity and expertise. Our extensive testing on four downstream tasks, notably including two innovative and practical tasks such as zero-shot question and answering and zero-shot virtual screening, demonstrate the efficacy of AMOLE in grasping the nuances of molecules and their textual descriptions.
@@ Appendix A Implementation Details
Text Encoder $f_{\text{text}}$. Following previous work Liu et al. (2023a), we use BERT architecture Devlin et al. (2018) as the text encoder $f_{\text{text}}$, and adapt the SciBERT Beltagy et al. (2019) checkpoint to initialize the model parameters  [^2] https://huggingface.co/allenai/scibert_scivocab_uncased. Text encoder contains a total 109,918,464 number of parameters.
Molecule Encoder $f_{\text{mol}}$. We use GIN Xu et al. (2018) architecture as a molecular encoder $f_{\text{mol}}$, which has been widely used as the backbone model in recent graph self-supervised learning works Hu et al. (2019); Liu et al. (2021). Additionally, we initialize the encoder parameters using the GraphMVP Liu et al. (2021) checkpoints provided by the original authors  [^3] https://huggingface.co/chao1224/MoleculeSTM/tree/main/pretrained_GraphMVP. Molecule encoder contains a total 1,885,206 number of parameters.
@@
Table 6: Hyperparameter specifications for AMOLE pretraining.
@@ Appendix B Baseline Methods
Single Encoder Models. For single encoder models, we mainly compare AMOLE with MolT5, BioT5, and KV-PLM. While T5-based models were not originally created for capturing the representations of molecules and text descriptions, we evaluate their performance by leveraging the hidden representations, aligning with approaches used in prior studies Seidl et al. (2023).
  • MolT5 Edwards et al. (2022) stands out as a trailblazer in the field of molecule language models. It introduces a novel approach of pre-training the model on an extensive dataset of unlabeled natural language texts and molecular strings (SMILES), using a denoising objective. Subsequently, the model undergoes fine-tuning for tasks such as molecule captioning and generation. For evaluation, we utilize the checkpoints provided by the authors, which are available in the Huggingface repository  [^4] https://huggingface.co/laituan245/molt5-large-caption2smiles.
  • BioT5 Pei et al. (2023) extends MolT5 by incorporating a diverse array of pre-training tasks, such as denoising molecule SELFIES, protein FASTA sequences, general text, wrapped text, as well as translating between bio-sequences and structured text descriptions. We also utilize checkpoints provided by the authors, which are available in the Huggingface repository  [^5] https://huggingface.co/QizhiPei/biot5-base-text2mol.
  • KV-PLM Zeng et al. (2022) introduces a pre-training approach for language models that incorporates masked language modeling on a specialized corpus featuring inserted SMILES strings. We utilize the checkpoint available on the author’s Github repository  [^6] https://github.com/thunlp/KV-PLM?tab=readme-ov-file.
@@
  • MoleculeSTM Liu et al. (2023a) suggests developing representations for molecules and texts through contrastive learning, where a molecule and its corresponding text are considered a positive pair, and all other combinations within the same batch are treated as negative pairs. Additionally, the authors introduce the extensive dataset, named PubChemSTM, which includes 250K distinct molecules and 281K molecule-text pairs. They introduce two distinct models, one utilizing SMILES representations of molecules as model input and the other as graph representations. For the SMILES model, they initially use the MegaMolBART Irwin et al. (2022) checkpoint, and for the graph model, the GraphMVP Liu et al. (2021) checkpoint. However, due to issues with the CUDA environment, we utilize ChemBERTa Chithrananda et al. (2020)  [^7] https://huggingface.co/seyonec/ChemBERTa-zinc-base-v1 as the initial checkpoint for the SMILES model training and GraphMVP  [^8] https://huggingface.co/chao1224/MoleculeSTM/tree/main/pretrained_GraphMVP for the graph model in MoleculeSTM.
  • MoMu Su et al. (2022) introduces a method of contrasting among molecules themselves with an additional loss function besides contrasting molecules and texts.
@@
  • Attribute Masking (AttrMask) Hu et al. (2019) introduces a technique of randomly masking the attributes of nodes or edges in the input and pre-training a GNN to predict these masked attributes.
  • Context Prediction (ContextPred) Hu et al. (2019) proposes to pre-train GNN to ensure nodes situated in analogous structural contexts are represented by proximate embeddings by using subgraphs to predict their surrounding graph structures.
  • GPT-GNN Hu et al. (2020) introduces a method for pre-training GNN through the generation of attributed graphs, which is achieved by dividing the generation process into two distinct phases: the generation of node attributes and the formation of edges between nodes.
  • InfoGraph Sun et al. (2019) proposes to maximize the mutual information between graph-level representation and node-level representation via contrastive learning.
  • GraphMVP Liu et al. (2021) introduces a hybrid training approach that merges generative and contrastive methodologies. In the contrastive setting, 2D molecular graphs and their corresponding 3D structures are considered positive pairs, while all other combinations are treated as negative pairs. On the generative side, the model aims to reconstruct the representation of a 3D structure from its 2D molecular graph counterpart and vice versa, facilitating a comprehensive understanding of molecular structures from both dimensions.
@@ Appendix C Datasets
Figure 4: (a) Histogram and (b) Boxplot on the number of descriptions per molecule.
Figure 5: Histogram on the number of words per description.
C.1 Pre-training
We pre-train AMOLE with the PubChem database, which is one of the most extensive public molecular databases available. PubChem database consists of multiple data sources including DrugBank, CTD, PharmGKB, and more. Please refer to the following URL for more details: https://pubchem.ncbi.nlm.nih.gov/sources/.
The PubChem database we used during training comprises a total of 299K unique molecules and 336K molecule-text pairs. During preprocessing, we consolidate each expertise into a unified description. Thus, each description for an individual molecule originates from a distinct database (expertise). On average, molecules are associated with 1.115 descriptions, with a maximum of 17 descriptions and a minimum of one. We provide a histogram and boxplot on the number of descriptions per molecule in Figure 4. As shown in Figure 4, it is evident that the majority of molecules have a singular description, indicating a lack of comprehensive expertise across various experts for numerous molecules. Moreover, each description in training data consists of 17.62 words on average, with a maximum of 874 words and a minimum of one. We also provide a histogram on the number of words per description in Figure 5.
C.2 Zero-Shot Cross-Modal Retrieval
Following previous work Liu et al. (2023a), we use the datasets extracted from the Description field, the Pharmacodynamics field, and the anatomical therapeutic chemical (ATC) field in the DrugBank database Wishart et al. (2018). Each field contains the following information:
@@ C.2 Zero-Shot Cross-Modal Retrieval
C.3 Zero-Shot Question and Answering
@@
Figure 6: Overall pipeline for generating and validating questions and answering datasets.
Figure 7: Prompts for (a) generating and (b) validating question and answer from textual description of molecule.
Figure 8: Failure cases in generating questions and answers.
Figure 9: Evaluation scheme for question and answering task.
@@
C.4 Molecular Property Prediction
For the molecular property prediction task, we use the MoleculeNet benchmark Wu et al. (2018) for evaluation, which is commonly used for evaluating the machine learning methods for molecular property prediction Liu et al. (2021). The MoleculeNet benchmark encompasses a diverse array of datasets, each characterized by distinct properties as follows:
@@ C.4 Molecular Property Prediction
C.5 Zero-Shot Virtual Screening
For the zero-shot virtual screening task, we utilize datasets from the Therapeutics Data Commons (TDC) Huang et al. (2021) [^10] https://tdcommons.ai/ and LIT-PCBA Tran-Nguyen et al. (2020) [^11] https://drugdesign.unistra.fr/LIT-PCBA/, containing the drugs that exhibit a range of desirable properties. Among the various datasets, we utilize the following datasets:
@@ C.5 Zero-Shot Virtual Screening
In addition to the previously mentioned datasets, we explore the virtual screening capabilities of the methods detailed in Appendix E.3 on the following datasets:
  • The oral bioavailability (Bioavailability) dataset encompasses 640 drugs, each labeled to indicate the extent to which the drug’s active ingredient is absorbed into the systemic circulation and becomes accessible at the intended site of action.
  • The blood-brain barrier (BBB) serves as a protective shield, preventing most foreign substances from entering. This dataset encompasses 1,975 drugs, each annotated with their ability to penetrate the BBB, posing a significant challenge in developing drugs for the central nervous system.
  • The human ether-a-go-go related gene (hERG) plays a vital role in regulating the heart’s rhythm. This dataset includes 648 drugs, each evaluated for their potential to block hERG, a condition that may result in significant adverse effects.
  • The HIV dataset contains 41,127 drugs and the label about its ability to inhibit HIV replication.
  • The SARS-CoV-2 dataset contains 1,480 drugs and the label about their activity against SARSCoV2.
@@ Appendix D Experimental Setups
D.1 Zero-Shot Cross-Modal Retrieval
In this section, we provide further details on experimental setups for the zero-shot cross-modal retrieval task. Following previous work Liu et al. (2023a), we evaluate the task performance in two distinctive settings 1) given molecule to retrieve the textual description, and 2) given texture description to retrieve the molecule. In each scenario, we conducted experiments with a range of options, specifically 4, 10, and 20 choices. Within these options, one is the matching counterpart, while the others are randomly selected from the dataset. Following this, the model performance is determined by its capacity to pinpoint the correct counterpart from the options provided, such as correctly matching the description to the given molecule or vice versa. For the evaluation, we conducted five separate experiments, each featuring a distinct random selection of options, and we present both the mean and standard deviation of these experiments.
D.2 Zero-Shot Question and Answering
@@
Dataset Prompt
HIA Human intestinal absorption (HIA)
Pgp Inhibition P-glycoprotein Inhibition
drug metabolism, and brain penetration, and its inhibition can seriously alter a drug’s bioavailability and safety
DILI Inducing liver injury
clavulanate isoniazid, and nonsteroidal anti-inflammatory drugs.
VDR Vitamin D receptor
Bioavailability Oral Bioavailability
The molecule is positive w.r.t. a property that is defined as ‘the rate and extent to which the
active ingredient or active moiety is absorbed from a drug product and becomes available at the site of action’
BBB Blood-Brain Barrier penetration
The molecule is able to penetrate the Blood-Brain Barrier, which is the protection layer that blocks most foreign drugs
as a membrane separating circulating blood and brain extracellular fluid, to deliver to the site of action.
hERG hERG Blocker
This molecule blocks the hERG, which is crucial for the coordination of the heart’s beating.
HIV Active against HIV virus
This molecule is active against HIV virus. These drugs typically possess one or more of the following properties:
Reverse Transcriptase Inhibitors, Protease Inhibitors, Integrase Inhibitors, Fusion Inhibitors, CCR5 Antagonists, Post-Attachment Inhibitors.
SARS-CoV-2 Active against SARS-CoV-2 virus
This molecule is active against SARS-CoV-2 virus. These drugs typically possess one or more of the following properties:
Viral Entry Inhibitors, RNA Polymerase Inhibitors, Protease Inhibitors, Immune Modulators, Interferon Inducers, Antibody-based Therapies.
Table 7: The abstractive and detailed prompts used for each dataset during the virtual screening task.
D.3 Molecular Property Prediction

### v3

@@ 5.5. Zero-Shot Virtual Screening
drug metabolism, and brain penetration, and its inhibition can seriously alter a drug’s bioavailability and safety.
@@
Figure 4. Hit rate (%) in zero-shot virtual screening task.
@@
Task Description. Virtual screening is a computational technique to search large libraries of compounds quickly to identify those structures most likely to have desired properties, emerging as a principal technique in the drug discovery process (Mehta et al., 2021). To examine MoLM’s real-world applicability, we propose a novel task named “Zero-shot Virtual Screening,” in which we assess the model’s proficiency in virtual screening drugs by supplying a textual description of a desired property. Specifically, given a dataset with binary labels, we first obtain the representation of textual description of the desired property from language model $f_{\text{text}}$. Then, we identify the top 100 molecules that are closest to the textual representation within the representation space of the dataset and calculate the hit rate to assess the model performance. Moreover, to further evaluate models’ robustness on the various textual descriptions, we conduct the experiment by providing two distinct descriptions for each property as shown in Table 6: one brief and abstract, and the other longer and more detailed, offering comprehensive information about the property. We provide details on each dataset in Appendix C.4.
@@
Empirical Results. In Figure 4, we have the following observations: 1) While previous studies have shown comparable results on various downstream tasks, their effectiveness in virtual screening leaves room for improvement. Notably, these models often yield results inferior to the mere random selection of molecules (i.e., gray dashed line in Figure 4), and their performance significantly fluctuates based on the specific textual description used. 2) However, AMOLE consistently performs the best in various datasets and different types of descriptions, further demonstrating its ability to learn a more qualified joint space between molecules and text. 3) One interesting observation is that, AMOLE demonstrates notably robust performance compared to the baseline methods, regardless of whether the description is abstract or detailed. This can be attributed to the expertise transfer module, which equips the model with the ability to deduce related information even when only an abstract level of detail is provided.
Figure 5. Sensitivity analysis on $\alpha$.
Sensitivity Analysis on $\mathbf{\alpha}$. In this section, we further investigate how the expertise reconstruction ($ER$) loss weighting factor, i.e., $\alpha$, affects the model robustness on virtual screening performance. In Figure 5, we observe that the model’s performance remains largely consistent when provided with a detailed and lengthy prompt, its performance on abstract and brief prompts significantly varies with the choice of $\alpha$. Specifically, the model’s effectiveness on abstract and concise prompts improves as $\alpha$ increases. This suggests that the expertise transfer module effectively allows the model to infer related information, thereby enhancing stability and performance, even when faced with brief and abstract prompts. In conclusion, AMOLE can consistently screen the qualified molecules given any textual description, highlighting the adaptability of AMOLE for real-world drug discovery process.
6. Conclusion and Future Works
In this paper, we introduce AMOLE, which tackles two distinct challenges in MoLM that set it apart from VLM: the limited availability of molecule-text paired data in terms of both quantity and expertise. To address these issues, we propose to selectively share the descriptions among the structurally similar molecules based on their structural similarity, and transfer the expertise between molecules by training the model to reconstruct one description from another. Our extensive evaluation of four downstream tasks, including two innovative and practical tasks such as zero-shot question and answering and zero-shot virtual screening, demonstrate the efficacy of AMOLE in grasping the nuances of molecules and their textual descriptions, highlighting its promising applicability for practical use in the drug discovery domain. On the other hand, while this paper suggests augmenting the molecule-text pairs by sharing descriptions among similar molecules, this method does not increase the total number of descriptions. As an interesting future direction, we suggest potential strategy to augment the absolute number of descriptions involves generating new descriptions based on the recent achievements of decoder-only models like GPT, and adapting the style of one expert to another.
Acknowledgement. This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2024-00335098), Ministry of Science and ICT (NRF-2022M3J6A1063021), and by the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2022-0-00077).
@@ Appendix A Implementation Details
Text Encoder $f_{\text{text}}$. Following previous work (Liu et al., 2023b), we use BERT architecture (Devlin et al., 2018) as the text encoder $f_{\text{text}}$, and adapt the SciBERT (Beltagy et al., 2019) checkpoint to initialize the model parameters  [^3] https://huggingface.co/allenai/scibert_scivocab_uncased. Text encoder contains a total 109,918,464 number of parameters.
Molecule Encoder $f_{\text{mol}}$. We use GIN (Xu et al., 2018) architecture as a molecular encoder $f_{\text{mol}}$, which has been widely used as the backbone model in recent graph self-supervised learning works (Hu et al., 2019; Liu et al., 2021). Additionally, we initialize the encoder parameters using the GraphMVP (Liu et al., 2021) checkpoints provided by the original authors  [^4] https://huggingface.co/chao1224/MoleculeSTM/tree/main/pretrained_GraphMVP. Molecule encoder contains a total 1,885,206 number of parameters.
@@
Table 7. Hyperparameter specifications for AMOLE pretraining.
@@ Appendix B Baseline Methods
Single Encoder Models. For single encoder models, we mainly compare AMOLE with MolT5, BioT5, and KV-PLM. While T5-based models were not originally created for capturing the representations of molecules and text descriptions, we evaluate their performance by leveraging the hidden representations, aligning with approaches used in prior studies (Seidl et al., 2023).
  • MolT5 (Edwards et al., 2022) stands out as a trailblazer in the field of molecule language models. It introduces a novel approach of pre-training the model on an extensive dataset of unlabeled natural language texts and molecular strings (SMILES), using a denoising objective. Subsequently, the model undergoes fine-tuning for tasks such as molecule captioning and generation. For evaluation, we utilize the checkpoints provided by the authors, which are available in the Huggingface repository  [^5] https://huggingface.co/laituan245/molt5-large-caption2smiles.
  • BioT5 (Pei et al., 2023) extends MolT5 by incorporating a diverse array of pre-training tasks, such as denoising molecule SELFIES, protein FASTA sequences, general text, wrapped text, as well as translating between bio-sequences and structured text descriptions. We also utilize checkpoints provided by the authors, which are available in the Huggingface repository  [^6] https://huggingface.co/QizhiPei/biot5-base-text2mol.
  • KV-PLM (Zeng et al., 2022) introduces a pre-training approach for language models that incorporates masked language modeling on a specialized corpus featuring inserted SMILES strings. We utilize the checkpoint available on the author’s Github repository  [^7] https://github.com/thunlp/KV-PLM?tab=readme-ov-file.
@@
  • MoleculeSTM (Liu et al., 2023b) suggests developing representations for molecules and texts through contrastive learning, where a molecule and its corresponding text are considered a positive pair, and all other combinations within the same batch are treated as negative pairs. Additionally, the authors introduce the extensive dataset, named PubChemSTM, which includes 250K distinct molecules and 281K molecule-text pairs. They introduce two distinct models, one utilizing SMILES representations of molecules as model input and the other as graph representations. For the SMILES model, they initially use the MegaMolBART (Irwin et al., 2022) checkpoint, and for the graph model, the GraphMVP (Liu et al., 2021) checkpoint. However, due to issues with the CUDA environment, we utilize ChemBERTa (Chithrananda et al., 2020)  [^8] https://huggingface.co/seyonec/ChemBERTa-zinc-base-v1 as the initial checkpoint for the SMILES model training and GraphMVP  [^9] https://huggingface.co/chao1224/MoleculeSTM/tree/main/pretrained_GraphMVP for the graph model in MoleculeSTM.
  • MoMu (Su et al., 2022) introduces a method of contrasting among molecules themselves with an additional loss function besides contrasting molecules and texts.
@@
  • Attribute Masking (AttrMask) (Hu et al., 2019) introduces a technique of randomly masking the attributes of nodes or edges in the input and pre-training a GNN to predict these masked attributes.
  • Context Prediction (ContextPred) (Hu et al., 2019) proposes to pre-train GNN to ensure nodes situated in analogous structural contexts are represented by proximate embeddings by using subgraphs to predict their surrounding graph structures.
  • GPT-GNN (Hu et al., 2020) introduces a method for pre-training GNN through the generation of attributed graphs, which is achieved by dividing the generation process into two distinct phases: the generation of node attributes and the formation of edges between nodes.
  • InfoGraph (Sun et al., 2019) proposes to maximize the mutual information between graph-level representation and node-level representation via contrastive learning.
  • GraphMVP (Liu et al., 2021) introduces a hybrid training approach that merges generative and contrastive methodologies. In the contrastive setting, 2D molecular graphs and their corresponding 3D structures are considered positive pairs, while all other combinations are treated as negative pairs. On the generative side, the model aims to reconstruct the representation of a 3D structure from its 2D molecular graph counterpart and vice versa, facilitating a comprehensive understanding of molecular structures from both dimensions.
@@ Appendix C Datasets
Figure 6. Overall pipeline for generating and validating questions and answering datasets.
Figure 7. Prompts for (a) generating and (b) validating question and answer from textual description of molecule.
Figure 8. Failure cases in generating questions and answers.
Figure 9. Evaluation scheme for question and answering task.
C.1. Zero-Shot Cross-Modal Retrieval
Following previous work (Liu et al., 2023b), we use the datasets extracted from the Description field, the Pharmacodynamics field, and the anatomical therapeutic chemical (ATC) field in the DrugBank database (Wishart et al., 2018). Each field contains the following information:
@@ C.1. Zero-Shot Cross-Modal Retrieval
C.2. Zero-Shot Question and Answering
@@
C.3. Molecular Property Prediction
For the molecular property prediction task in Section 5.4, we use the MoleculeNet benchmark (Wu et al., 2018) for evaluation, which is commonly used for evaluating the machine learning methods for molecular property prediction (Liu et al., 2021). The MoleculeNet benchmark encompasses a diverse datasets, each characterized by distinct properties as follows:
@@ C.3. Molecular Property Prediction
C.4. Zero-Shot Virtual Screening
For the zero-shot virtual screening task, we utilize datasets from the Therapeutics Data Commons (TDC) (Huang et al., 2021) and LIT-PCBA (Tran-Nguyen et al., 2020), containing the drugs that exhibit a range of desirable properties. Among the various datasets, we utilize the following datasets:
@@ Appendix D Experimental Setups
D.1. Zero-Shot Cross-Modal Retrieval
In this section, we provide further details on experimental setups for the zero-shot cross-modal retrieval task. Following previous work (Liu et al., 2023b), we evaluate the task performance in two distinctive settings 1) given molecule to retrieve the textual description, and 2) given texture description to retrieve the molecule. In each scenario, we conducted experiments with a range of options, specifically 4, 10, and 20 choices. Within these options, one is the matching counterpart, while the others are randomly selected from the dataset. Following this, the model performance is determined by its capacity to pinpoint the correct counterpart from the options provided, such as correctly matching the description to the given molecule or vice versa. For the evaluation, we conducted five separate experiments, each featuring a distinct random selection of options, and we present both the mean and standard deviation of these experiments.
D.2. Zero-Shot Question and Answering
@@
D.3. Molecular Property Prediction